OpenStudy (anonymous):

If f(x) = -1/3 x + 6 and g(x) = -3/4 x +9 find (f o g)(x)

OpenStudy (anonymous):

(f o g)(x)=-1/3 (-3/4 x +9) + 6

OpenStudy (anonymous):

simplify accordingly

OpenStudy (anonymous):

These are the answer choices A. (f o g)(x) = 1/4x^2 - 27/2x+54 B. (f o g)(x) = -1/4x+9 C. (f o g)(x) = 1/4 x + 3

OpenStudy (precal):

you can finish the problem. We can help you as you do it or check your solution